# Audience Impressions of Narrative Structures and Personal Language Style in Science Communication on Social Media

### v2

@@ Abstract.
Science communication increases public interest in science by educating, engaging, and encouraging everyday people to participate in the sciences. But traditional science communication is often too formal and inaccessible for general audiences. However, there is a growing trend on social media to make it more approachable using three techniques: relatable examples to make explanations concrete, step-by-step walkthroughs to improve understanding, and personal language to drive engagement. These techniques are flashy and often garner more engagement from social media users, but the effectiveness of these techniques in actually explaining the science is unknown. Furthermore, many scientists struggle with adopting these science communication strategies for social media, fearing it might undermine their authority. We conduct a reader study to understand how these science communication techniques on social media affect readers’ understanding and engagement of the science. We found that while most readers prefer these techniques, they had diverse preferences for when and where these techniques are used. With these findings, we conducted a writer study to understand how scientists’ varying comfort levels with these strategies can be supported by presenting different structure and style options. We found that the side-by-side comparison of options helped writers make editorial decisions. Instead of adhering to one direction of science communication, writers explored a continuum of options which helped them identify which communication strategies they wanted to implement.
@@ 1. Introduction
Science communication increases public interest in science by educating, engaging, and encouraging everyday people to participate in the sciences (of Sciences et al., 2017; Burns et al., 2003). With the rise of social media, science communication has transitioned away from traditional scholarly articles and towards more informal platforms such as blogs and social media (Hou and Wang, 2017; Williams et al., 2022; Brüggemann et al., 2020a). There is a growing need for scientists to participate in these public forums to bridge the gap between science and everyday people (Pearson, 2001; Powell and Lee Kleinman, 2008).
@@
On social media, science communication strategies emphasize making science engaging and understandable to non-technical audiences and adapt common practices in science communication literature for a social media audience (Gilbert, 2020). A study of science communication on Twitter found these strategies: 1) use of a relatable example – such as using the iridescence of bubbles to motivate the topic “Thin Film Interference” in physics, 2) use of a step-by-step walkthrough that uses 10-12 tweets which use an example to explain every step, and 3) use of personal language – the writer talking about themselves and their experience, often in a personal and conversational way, “Crazy, right? WTH was I thinking ?!?” (Gero et al., 2021a).
But previous work has not evaluated whether these science communication techniques on social media (example, walkthrough, and personal language) are actually helpful in helping everyday people learn about science. As such, we conduct a study of non-technical readers’ rating of STEM explanations with and without each of these three social media explanation techniques to understand how each factor impacts making science explanations engaging and understandable. To facilitate fair comparisons between each condition, we use an LLM to generate science explanations for 15 topics: 5 STEM fields, each with 3 topics at different levels of complexity. Readers evaluated them for understanding and engagement. Overall, readers preferred explanations with examples over no examples and personal language over no personal language. However, readers were split between narratives with and without walkthroughs; certain topics were more suitable for narratives without a walkthrough. Even for techniques they did prefer, readers’ personal experience complicate why some preferred narratives without a given technique. There is no one-size-fits-all approach to science communication on social media.
As such, it’s important to support STEM experts in navigating diverse audience preferences and help them engage with science communication on social media. Many STEM experts are eager to communicate to the public (Della Giusta et al., 2021). But many scientists are not trained in how to translate complex technical topics for an everyday audience and struggle to motivate and explain topics to people beyond their expertise (Williams et al., 2022). Furthermore, studies find that scientists are hesitant in adopting science communication strategies for social media because inaccurate interpretations or misrepresentations of their work could affect public understanding and their own reputations (Loroño-Leturiondo and Davies, 2018; Gero et al., 2021b). Scientists often feel conflicted over their personal and professional identities when engaging in science communication on social media (Koivumäki et al., 2020).
We explore how to accommodate different writers’ comfort levels when it comes to adopting science communication strategies for social media. We ran an in-depth 2-hour study of writers (n=10) who were given three initial draft options with different narrative structures (One Example, No Example, and Many Examples) and selected one draft to move forwards with. Then, writers saw the selected narrative presented With Personal Language and Without Personal Language. We found that while many writers maintained their beliefs about how science should be communicated, seeing options side-by-side helped them recognize effective and ineffective science communication techniques, ultimately helping them make editorial decisions regarding what to cut, edit, and merge. Furthermore, when presented with science explanations with and without personal language, scientists often chose to merge lines of dialog between the different conditions. Overall, this shows there is continuous design space for structure and style options in public science communication. Writers can be aided by seeing points in the design space and finding a place they are comfortable with on the continuum.
We conclude with a discussion of how to mediate between audiences’ needs for understanding and engagement and accommodating writers’ hesitancy around loss of authority and oversimplification of science.
@@ 2. Background on Tweetorials
Tweetorials are a form of social media science communication on Twitter, defined as a chronological series of tweets that explains a science topic (Symplur, 2019; Brüggemann et al., 2020b). According to Breu and Berstein, Tweetorials emerged from the medical community to continue education for other medical professionals, often containing medical jargon and not intended for everyday audiences. As the form became popularized, other communities on Twitter began to adopt the format (Breu, 2019; Symplur, 2019; Brüggemann et al., 2020b). As a result, Breu’s definition of a Tweetorial, “a collection of threaded tweets aimed at teaching users who engage with them,” can be applied to various domains such as biology, computer science, and economics.
Gero et. al. identified and compared key features of Tweetorials to traditional science communication strategies (Gero et al., 2021b). The authors identified 3 main structural components of Tweetorials to be the hook, body, and conclusion. Previous research has explored strategies for writing engaging hooks (Long et al., 2024, 2023). Thus, we focus on specific strategies that appear within the body of a Tweetorial.
@@ 2.1. Key Strategies of the Tweetorial Body
According to Gero et. al., the body of a Tweetorial is the most varied in length and types of detail used to explain a given topic (Gero et al., 2021b). The researchers found that Tweetorials contain specific techniques such as the use of an example, a step-by-step walkthrough structure, and personal language.
@@ 2.1.1. Example (E)
Like traditional science communication, Tweetorials often contain an example that uses familiar or simpler concepts to explain the main idea (Gero et al., 2021b). We define this as an example (E) technique. Components of the example are the use case and the scenario. The use case is a general application of the scientific topic and the scenario is a specific situation that describes how the science topic was applied. For example, one Tweetorial uses the example of fingertips getting wrinkly in the bath to explain water immersion wrinkling. [^1] http://language-play.com/tech-tweets/tweetorial/4 For this Tweetorial, the use case is when fingers get wrinkly in water. The scenario is a parent bathing their child. We use the use case and scenario for more finegrain control over the LLM-generated narratives in Section 4.2.
@@ 3.1. HCI for Science Communication
Science communication on social media is informal and often told from a personal perspective to engage a broader audience (Gero et al., 2021b). Previous work has compiled recommendations for effective science communication on social media based on how likely using these strategies will increase engagement: comments, likes, and shares (Zhang and Lu, 2022; Bik et al., 2015; Fontaine et al., 2019). But these works do not consider whether these techniques are actually effective for readers to understand the science. Our work evaluates whether these different science communication strategies on Twitter are effective in helping readers engage with and understand the science.
@@ 4.1. Research Questions
Tweetorials typically include three techniques example (E), walkthrough (W), and personal language (P) (Gero et al., 2021b). We want to understand how each of these techniques affects readers’ ratings of science communication on social media. Given that most published Tweetorials contain these three techniques, we hypothesize that science explanations that contain all three features, EWP, will have the highest reader preference rating. To evaluate the effect of the 3 different features, we compare narratives with all three features (EWP) to narratives with one of the features removed (EW, EP, WP). We investigate the following hypotheses in a survey study with readers:
@@ 5. Results: Reader Study
5.1. Quantitative Findings on Survey Data
Overall, we collected 105 ratings on 75 different science explanations. This consisted of 35 different annotators who evaluated each science explanation on 4 different Likert-scale questions. We conduct a quantitative analysis of the survey data followed by a qualitative analysis of individual reader’s preferences for science communication to contextualize nuances to the quantitative data.
5.1.1. Usage of Example Preferences
@@
Table 1. GLMM Results for H1:Examples
@@
5.1.2. Step-by-Step Walkthrough Preferences
To evaluate H2:Walkthrough, we compare experimental condition EWP-NoFewShot to baseline condition EP-NoFewShot. We found that readers had a slight preference for explanations with a walkthrough compared to without a walkthrough, with a considerable variance on the fields of study. Table 2 summarizes the findings from a GLMM analysis. The experimental condition, EWP-NoFewShot received a score of 15.965 which was not statistically significant compared to 14.729, the baseline score for EP-NoFewShot (p = 0.057). As such, there is no significant difference in how readers rated explanations with and without walkthroughs. The random effects analysis shows a minimal group variance of 0.002, indicating little variability among participants. However, the field variance is considerable at 2.566, suggesting that participants prefer having a walkthrough for some STEM topics but prefer having no walkthrough for other STEM topics. Our qualitative findings (Section 5.2) provide further explanations on why some participants are divided in their preferences for walkthrough.
Table 2. GLMM Results for H2:Walkthrough
@@ 5.1.2. Step-by-Step Walkthrough Preferences
5.1.3. Personal Language Preferences
To evaluate H3:Personal Language, we compare experimental condition EWP to baseline condition EW. We found that participants prefer reading explanations with personal language over without personal language. Table 3 summarizes the findings from a GLMM analysis. The experimental condition, EWP, received a statistically significantly higher score of 15.883 compared to 13.973 for the baseline condition, EW (p ¡ 0.0001). The effect size is 1.910 which means that readers rate explanations with personal language almost 2 points higher than explanations without personal language. Additionally, there is a group variance of 1.392 indicating a moderate level of variability among participants and a fields variance of 1.361 showing a moderate variance among STEM topics.
Table 3. GLMM Results for H3:Personal Language
@@ 5.1.3. Personal Language Preferences
5.2. Qualitative Findings on Reader’s Preferences
Overall, participants preferred science explanations with an example over no example and with personal language over no personal language. However, participants were split on the their preference for narratives with and without walkthroughs. We use semi-structured interviews to gain further insights into the reasons why participants might prefer narratives with or without an example, walkthrough, and personal language.
5.2.1. Most Readers Preferred Examples
Overall, readers reported that having an example was helpful. 4 of 8 readers stated the example helped them understand the importance of a topic (P2, P4, P7, P8). 5 of 8 found that examples helped them reflect and understand their own experiences better (P1, P2, P3, P4, and P7). When reading a science explanation about the computer science topic of depth-first search, P2 remarked that learning how the algorithm can be applied to navigating a maze helped her engage with the topic: “I don’t really care about computer science algorithms, but I do care about how this applies to my own life.”
However, not all readers needed an example to help them understand the topic. 3 of 8 participants mentioned how the example felt unnecessary and detracted from the content of the explanation (P1, P2, P8). When reading about depth-first search, P1 mentioned how the given example of navigating through the Botanical Garden felt unnecessary: “The concept in and of itself is interesting. I would have just read about that. I don’t really need any more context.” For P8, when learning about thin film interface through the example of a child playing with bubbles, he remarked that “[the example] doesn’t really pertain to me in any way,” demonstrating how certain examples might not resonate with particular audiences.
3 of the 8 participants mentioned how they would reference their own experiences to help ground the technical explanation (P1, P2, P4). When P2 was reading about thin film interference without an example, she used her own experience in working with glass to ground her understanding of the topic: “[the topic] was really relevant to my life and what I’m already doing.” For P4, when the given example of watching a horror film was used to explain Walker’s Action Decrement Theory, she mentioned how even though the example was unrelatable, she referenced her own experiences to find an example that fit the same context. This shows that readers use their own experiences to contextualize the science for explanations that do not include an example or that include an example unrelatable to the reader.
5.2.2. Readers had No Preference for Narratives with Walkthroughs
There was no statistical significance in our results comparing reader preferences for narratives with and without walkthroughs. In this section we will investigate why some readers prefer walkthroughs or no walkthroughs.
Some readers reported that the walkthrough helped establish a structure for the explanation (P2, P3, P8), provide evenly paced information (P1, P3, P6), and helped them follow through with the science (P4, P6, P7). P3 appreciated the walkthrough for the topic of Walker’s Action Decrement Theory in Psychology because it helped them understand how their heightened emotional state when watching a horror movie might affect their memory. By walking through 3 different stages of the narrator’s emotional journey during this event and how it affects their memory, P3 mentioned how the narrative helped “set the stage” and “allows you to follow through with [the science explanation].” The walkthrough provided a scaffold to support the reader in processing the information as a sequence of events.
But not all readers preferred narratives with a walkthrough. 3 participants (P2, P3, P7) reported that the walkthrough structure made the explanation feel overly explanatory or repetitive. For the topic of Walker’s Action Decrement Theory, P2 said that the walkthrough narrative was “just overly explanatory for a concept that is very intuitive,” demonstrating how for certain topics a walkthrough might not be necessary.
4 participants (P2, P4, P5, P6) preferred science explanations without a walkthrough but with many examples because they provided multiple different angles to view a topic in a condensed space. P4 stated that the explanation without a walkthrough broke down the topic of the curtain wall system into smaller, self-contained chunks of information that made reading the explanation “less intimidating.” For example, one paragraph of the explanation focused specifically on the aspect of temperature control, while the subsequent paragraph explored the structural construction of the curtain wall. P4 said that having the information broken down in this way made it more “digestible” to learn about the topic. For P6, the “list of facts” structure of explanation without a walkthrough helped them gain a broad overview of the topic. This illustrates that for some topics, readers might not want an in-depth walkthrough of the science and might prefer getting a high-level overview of how the science works.
5.2.3. Personal Language Sometimes Distracts from the Science
Overall, readers preferred reading explanations with personal language. Participants reported that the personal language helped establish a writer/reader connection (P1, P3, P5). But not all readers preferred narratives with personal language. 2 readers preferred reading science explanations with no personal language because they believed personal language was unnecessary (P1, P8): “[the explanation is] full of fluff coming from a personal perspective that I didn’t really care about” (P1). P8 mentioned that he did not need personal language for engagement when reading about science, additionally, he hypothesized that other readers “[might] need the personal aspects to get them to read something that they wouldn’t otherwise be interested in.” Depending on a reader’s inclination towards science, personal language may or may not be necessary to help them engage with these science narratives.

### v3

# “I never would have thought to say this”: Example-Based Exploration to Balance Scientists’ Writing Preferences with Public Science Communication Strategies

@@ Abstract.
Public-facing science communication is important in garnering interest, engagement, and trust in science. Social media platforms provide scientists with opportunities to reach broader audiences, yet many resist adopting social media writing strategies because the strategies conflict with traditional science writing norms and personal preferences. To address this gap, we first evaluate readers’ preferences for strategies such as examples, walkthroughs, and personal language. While many readers enjoyed science narratives that used these strategies, their effectiveness was nuanced and context-dependent, varying by topic and individual preference. Building on these findings, we design a system that uses contrastive examples to help scientists adopt and integrate these social media science writing strategies. In a user study with scientists, we found that presenting contrastive examples helped writers critically evaluate different narrative options, balance competing goals, and gain confidence in adapting social media writing strategies to fit both their topic and audience.
@@ 1. Introduction
Scientists go through years of training to learn the structure and style of writing research papers as a way for other scientists in their field to learn about their work (Pearson, 2001; Powell and Lee Kleinman, 2008). But sharing science beyond these formal research communities is important in supporting public interest in science (of Sciences et al., 2017; Burns et al., 2003). Science communication on social media platforms, like X, has emerged as a straightforward and accessible way for scientists to communicate with the public (Hou and Wang, 2017; Williams et al., 2022; Brüggemann et al., 2020b). To do so, scientists must learn a different set of writing structures and styles to use when communicating science to an everyday audience on social media (Gilbert, 2020; Della Giusta et al., 2021).
Prior work has identified specific strategies of successful science posts around how to connect with an everyday audience via relatable examples, a step-by-step walkthrough, and personal language (Gero et al., 2021b). But this type of writing is often at odds with how scientists were trained to write about science. Prior work has found that scientists struggle to negotiate their identities as scientists with the expectations of social media (Loroño-Leturiondo and Davies, 2018). For example, scientists might question whether using personal language will detract from their credibility as a scientist (Gero et al., 2021a; Koivumäki et al., 2020).
@@
In order to support scientists in engaging with this type of science communication, we need to understand readers’ preferences of these different writing strategies to help scientists feel more comfortable adopting social media writing strategies. To do so, we evaluate readers’ perspectives on these different social media science communication strategies: 1) use of a relatable example – such as using the iridescence of bubbles to motivate the topic “Thin Film Interference” in physics, 2) use of a step-by-step walkthrough that uses 10-12 tweets which use an example to explain every step, and 3) use of personal language – the writer talking about themselves and their experience, often in a personal and conversational way, “Crazy, right? WTH was I thinking ?!?” (Gero et al., 2021b). Prior research shows that teaching writers about these strategies doesn’t help them incorporate them into their writing for social media (Gero et al., 2021b). Using the findings from the reader study, we design a system that provides scientists with different examples of the science communication strategies to help scientists balance their own science writing background with these social media writing strategies.
Overall, readers did prefer the use of a relatable example and personal language, but were split for the use of a walkthrough. While many readers enjoyed the use of personal language to make the science narratives more engaging, there were readers who found personal language to be distracting to the science. Similarly, while many readers preferred examples, some readers found that the use of an example often added unnecessary details to explaining science or isolated readers because the example used was unrelatable. Because of readers’ differing preferences, these writing strategies do not work for all readers and their effectiveness varies based on the topic.
Given the diverse and contextual preferences that readers had for these science communication strategies, static writing guidelines do not accommodate these preferences. We design a system that presents scientists a set of options for how to (1) structure and (2) style their writing. We find that presenting each dimension of social media writing strategies through different options helps scientists balance their science background and readers’ preferences. Presenting different structure and style options allows writers to choose the degree to which they employ a given social media strategy, turning static writing strategies into a spectrum of options for writers to adopt according to their own preferences.
We find that showing scientists options allows them to more easily adopt these science communication strategies while still maintaining their own preferences and beliefs about how science should be communicated. Instead of presenting science communication as an immutable set of guidelines, we find that presenting writers with different structure and style options can support scientists in making more informed and intentional decisions about when and how to integrate these social media writing strategies.
@@ 2. Background on Tweetorials
Tweetorials are a form of social media science communication on Twitter, defined as a chronological series of tweets that explains a science topic (Symplur, 2019; Brüggemann et al., 2020a). According to Breu and Berstein, Tweetorials emerged from the medical community to continue education for other medical professionals, often containing medical jargon and not intended for everyday audiences. As the form became popularized, other communities on Twitter began to adopt the format (Breu, 2019; Symplur, 2019; Brüggemann et al., 2020a). As a result, Breu’s definition of a Tweetorial, “a collection of threaded tweets aimed at teaching users who engage with them,” can be applied to various domains such as biology, computer science, and economics.
Gero et. al. identified and compared key features of Tweetorials to traditional science communication strategies (Gero et al., 2021a). The authors identified 3 main structural components of Tweetorials to be the hook, body, and conclusion. Previous research has explored strategies for writing engaging hooks (Long et al., 2024, 2023). Thus, we focus on specific strategies that appear within the body of a Tweetorial.
@@ 2.1. Key Strategies of the Tweetorial Body
According to Gero et. al., the body of a Tweetorial is the most varied in length and types of detail used to explain a given topic (Gero et al., 2021a). The researchers found that Tweetorials contain specific techniques such as the use of an example, a step-by-step walkthrough structure, and personal language.
@@ 2.1.1. Example (E)
Like traditional science communication, Tweetorials often contain an example that uses familiar or simpler concepts to explain the main idea (Gero et al., 2021a). We define this as an example (E) technique. Components of the example are the use case and the scenario. The use case is a general application of the scientific topic and the scenario is a specific situation that describes how the science topic was applied. For example, one Tweetorial uses the example of fingertips getting wrinkly in the bath to explain water immersion wrinkling. [^1] http://language-play.com/tech-tweets/tweetorial/4 For this Tweetorial, the use case is when fingers get wrinkly in water. The scenario is a parent bathing their child. We use the use case and scenario for more finegrain control over the LLM-generated narratives in Section 4.2.
@@ 3.1. HCI for Science Communication
Science communication on social media is informal and often told from a personal perspective to engage a broader audience (Gero et al., 2021a). Previous work has compiled recommendations for effective science communication on social media based on how likely it is that using these strategies will increase engagement: comments, likes, and shares (Zhang and Lu, 2022; Bik et al., 2015; Fontaine et al., 2019). But these works do not consider whether these techniques are actually effective for readers to understand the science. Our work evaluates whether these different science communication strategies on Twitter are effective in helping readers engage with and understand the science.
@@ 4.1. Research Questions
Tweetorials typically include three techniques example (E), walkthrough (W), and personal language (P) (Gero et al., 2021a). We want to understand how each of these techniques affects readers’ ratings of science communication on social media. Given that most published Tweetorials contain these three techniques, we hypothesize that science explanations that contain all three features, EWP, will have the highest reader preference rating. To evaluate the effect of the 3 different features, we compare narratives with all three features (EWP) to narratives with one of the features removed (EW, EP, WP). We investigate the following hypotheses in a survey study with readers:
@@ 5. Results: Reader Study
Overall, we collected 105 ratings on 75 different science explanations. This consisted of 35 different annotators who evaluated each science explanation on 4 different Likert-scale questions. We conduct a quantitative analysis of the survey data followed by a qualitative analysis of interview transcripts (n=8) for individual reader’s preferences for science communication to contextualize nuances to the quantitative data.
5.1. Usage of Example Preferences
@@
From our qualitative interviews with readers, we found that overall, readers reported that having an example was helpful. 4 of 8 readers stated the example helped them understand the importance of a topic (P2, P4, P7, P8). 5 of 8 found that examples helped them reflect and understand their own experiences better (P1, P2, P3, P4, and P7). However, not all readers needed an example to help them understand the topic. 3 of 8 participants mentioned how the example felt unnecessary and detracted from the content of the explanation (P1, P2, P8). When reading about depth-first search, P1 mentioned how the given example of navigating through the Botanical Garden felt unnecessary: “The concept in and of itself is interesting. I would have just read about that. I don’t really need any more context.” For P8, when learning about thin film interface through the example of a child playing with bubbles, he remarked that “[the example] doesn’t really pertain to me in any way,” demonstrating how certain examples might not resonate with particular audiences.
These results demonstrate that while the use of examples generally help in providing additional context and grounding for a science narrative, the effectiveness of an example depends on the science topic and the readers’ personal preferences. Scientists do not need to avoid examples out of fear of distracting or alienating readers. Instead, the use of examples in science narratives is situational and scientists should have the agency to compare how different narratives with and without an example might resonate with readers depending on the topic and audience.
@@
Table 1. GLMM Results for H1:Examples show that readers demonstrate a statistically significant preference for science narratives with an example over narratives without an example.
5.2. Step-by-Step Walkthrough Preferences
To evaluate H2:Walkthrough, we compare experimental condition EWP-NoFewShot to baseline condition EP-NoFewShot. We found that readers had a slight preference for explanations with a walkthrough compared to without a walkthrough, with a considerable variance on the fields of study.
Table 2 summarizes the findings from a GLMM analysis. The experimental condition, EWP-NoFewShot received a score of 15.965 which was not statistically significant compared to 14.729, the baseline score for EP-NoFewShot (p = 0.057). As such, there is no significant difference in how readers rated explanations with and without walkthroughs. The random effects analysis shows a minimal group variance of 0.002, indicating little variability among participants. However, the field variance is considerable at 2.566, suggesting that participants prefer having a walkthrough for some STEM topics but prefer having no walkthrough for other STEM topics.
From our qualitative findings, we find that the use of a walkthrough helped helped establish a structure for the explanation (P2, P3, P8), provide evenly paced information (P1, P3, P6), and helped them follow through with the science (P4, P6, P7). But for 3 readers, the use of a walkthrough sometimes made the explanation feel overly explanatory or repetitive (P2, P3, P7). For the topic of Walker’s Action Decrement Theory, P2 said that the walkthrough narrative was “just overly explanatory for a concept that is very intuitive,” demonstrating how for certain topics a walkthrough might not be necessary. 4 participants (P2, P4, P5, P6) preferred science explanations without a walkthrough because they provided multiple different perspectives to understand the topic. P4 stated that the explanation without a walkthrough broke down the topic of the curtain wall system into smaller, self-contained chunks of information that made reading the explanation “less intimidating.”
These results demonstrate that while a step-by-step walkthrough was helpful in explaining the science, but for certain STEM topics the walkthrough might not be necessary. Scientists should use their own best judgment in evaluating when added detail supports clarity and when a more streamlined explanation may be sufficient for the topic at hand.
@@ 5.2. Step-by-Step Walkthrough Preferences
Table 2. GLMM Results for H2:Walkthrough shows that there is no significant difference between readers’ preferences for science narratives with or without walkthroughs.
5.3. Personal Language Preferences
To evaluate H3:Personal Language, we compare experimental condition EWP to baseline condition EW. We found that participants prefer reading explanations with personal language over without personal language.
Table 3 summarizes the findings from a GLMM analysis. The experimental condition, EWP, received a statistically significantly higher score of 15.883 compared to 13.973 for the baseline condition, EW (p ¡ 0.0001). The effect size is 1.910 which means that readers rate explanations with personal language almost 2 points higher than explanations without personal language. Additionally, there is a group variance of 1.392 indicating a moderate level of variability among participants and a fields variance of 1.361 showing a moderate variance among STEM topics.
From the qualitative findings, we see that science narratives with personal language helped establish a writer/reader connection (P1, P3, P5) and helped make science topics more approachable (P2, P3, P6, P7). But not all readers preferred narratives with personal language. 2 readers preferred reading science explanations with no personal language because they believed personal language was unnecessary (P1, P8): “[the explanation is] full of fluff coming from a personal perspective that I didn’t really care about” (P1). P8 mentioned that he did not need personal language for engagement when reading about science, though he hypothesized that other readers “[might] need the personal aspects to get them to read something that they wouldn’t otherwise be interested in.” Depending on a reader’s inclination towards science, personal language may or may not be necessary to help them engage with these science narratives. Even for readers who generally do prefer science narratives with personal language, there’s a balance between narratives with just enough and too much personal language (P2, P7). P2 compared two narratives with personal language saying how for one example “the language was not distracting” while the other example “had a ton of hashtags and full of like, fluff.”
While personal language is helpful in making science more engaging, the sentiment is not shared by all readers. Even for readers who generally enjoy personal language in science narratives, there’s a balance of how much personal language should be incorporated before it detracts from the science. These findings highlight that preferences are not binary; rather, they fall along a spectrum, where different readers find varying degrees of personal language to be effective for a science narrative.
@@ 5.3. Personal Language Preferences
Table 3. GLMM Results for H3:Personal Language show that there is a statistically significant difference in readers preference for science narratives that contain personal language over narratives without personal language.